OpenStudy (anonymous):

Given the functions m(x) = 4x − 11 and n(x) = x − 10, solve m[n(x)] and select the correct answer below.

OpenStudy (anonymous):

m[n(x)] = 4x − 51 m[n(x)] = 4x − 29 m[n(x)] = 4x^2− 51 m[n(x)] = 4x^2 − 29

OpenStudy (anonymous):

@ali2x2 @Majesty69 @sammixboo

OpenStudy (campbell_st):

you are being asked to substitute n(x) into m(x) so m(n(x)) = 4( x - 10) - 11 now simplify

OpenStudy (anonymous):

okay so =4x-51

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

so A

OpenStudy (campbell_st):

that's correct